Assume that interest rate parity holds. In the spot market 1 Japanese yen = $0.01392, while in the 90-day forward market 1 Japanese yen = $0.01395. In Japan, 90-day risk-free securities yield 2.5%. What is the yield on 90-day risk-free securities in the United States? Round your answer to two decimal places. Do not round intermediate calculations.

%
Answer #1

ANSWER

As per interest rate parity theory:

Forward rate = spot rate *[1+ interest rate in foreign country]/ [1+ interest rate in home country]

Forward rate = 0.01395

Spot rate =0.01392

Interest rate in home country =2.5% = 0.025

Let x be interest rate in foreign country i.e.: United States

0.01395= 0.01392*[1+x]/[1+0.025]

0.01395= 0.01392*[1+x]/1.025

0.01395*1.025 = 0.01392*[1+x]

0.01429875 = 0.01392*[1+x]

[1+x] = 0.01429875 /0.01392= 1.02720905172

X= 1.02720905172– 1 = 0.02720905172

Rounding the answer = 2.72%

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************